2. QUESTION:

A fair coin is tossed, and a fair die is thrown. Write down sample spaces for

(a) the toss of the coin;

(b) the throw of the die;

(c) the combination of these experiments.

Let A be the event that a head is tossed, and B be the event that an odd number is thrown. Directly from the sample space, calculate P(A ∩ B) and P(A ∪ B).


1
Expert's answer
2022-06-27T06:30:50-0400

(a) "\{HEAD,TAIL\}"

(b) "\{1,2,3,4,5,6\}"

(c)

"\{(HEAD,1), (HEAD,2), (HEAD,3),\\ (HEAD,4), (HEAD,5), (HEAD,6),\\
 (TAIL,1), (TAIL,2), (TAIL,3),\\ (TAIL,4), (TAIL,5), (TAIL,6)\}"

The sample space of combination of the experiments has "12" elements with equal probabilities "\frac{1}{12}."

"A\cap B=\{(HEAD,1), (HEAD,3), (HEAD,5)\}",

"|A\cap B|=3", hence "P(A\cap B)=3\cdot \frac{1}{12}=\frac{1}{4}."

"A\cup B=\{(HEAD,1), (HEAD,2), (HEAD,3),\\ (HEAD,4), (HEAD,5), (HEAD,6),\\
(TAIL,1), (TAIL,3), (TAIL,5)\},"

"|A\cup B|=9," hence "P(A\cup B)=9\cdot \frac{1}{12}=\frac{3}{4}."

Answer: "P(A\cap B)=\frac{1}{4}, P(A\cup B)=3\cdot \frac{1}{12}=\frac{3}{4}."